OpenStudy (anonymous):

"m1" =2, "m2"= -1/2 are the two lines perpendicular?

OpenStudy (anonymous):

two lines are perpendicular only if m1*m2=-1

OpenStudy (anonymous):

Is 2*(-1/2)=-1 ?

OpenStudy (anonymous):

u have to check if m1*m2 is equal to -1 or not

OpenStudy (anonymous):

no its not...... @sauravshakya

OpenStudy (anonymous):

why not??? 2*(-1/2)=-1

OpenStudy (anonymous):

it is equal to -1

OpenStudy (anonymous):